OpenStudy (anonymous):

For : tan (Ө - 45 degrees) + tan ( Ө + 45 degrees) = 2 tan2Ө, how do you apply the identity of : cot Ө = tan (90 degrees - Ө ) ? in order to verify the equation.

OpenStudy (anonymous):

try expanding them like tan(a-b) = ........... and simplify.........that shud do it

OpenStudy (anonymous):

\[\tan(A -B) = \frac{tanA - tanB}{1 + tanA.tanB}\] \[\tan(A +B) = \frac{tanA + tanB}{1 - tanA.tanB}\] \[\frac{2tanA}{1 - \tan^2A} = \tan2A\] Use these identities to verify...